OpenStudy (anonymous):

Question: 4___4___4___4 = 20 Use +, -, /, * To Solve It. It’s A Challenge.

OpenStudy (kinggeorge):

\((4/4 +4)*4\)

OpenStudy (anonymous):

where is ( - ) sign ?

OpenStudy (kinggeorge):

You can't use all 4 signs. There's only room for 3.

OpenStudy (anonymous):

no King George I bet one of the fours are a negative

OpenStudy (anonymous):

which leaves the other 3

OpenStudy (anonymous):

Well, it means there is no answer for this ?

OpenStudy (kinggeorge):

If we're going that route, \[\left(\frac{-4}{-4}+4\right)*4\]

OpenStudy (anonymous):

VICTORY! :D

OpenStudy (kinggeorge):

Alternatively, \[(-4/4-4)*(-4)\]

OpenStudy (anonymous):

wouldn't that be -12?

OpenStudy (kinggeorge):

\[-4/4=-1\]\[-1-4=-5\]\[-5*-4=20\]

OpenStudy (anonymous):

oh! I did the second part backwad, my bad

OpenStudy (anonymous):

*backwards

OpenStudy (anonymous):

yes, only changing signs of numbers we can solve it. Otherwise it is not possible, right?

OpenStudy (kinggeorge):

In my opinion, the puzzle isn't asking you to use all 4 symbols. Rather, use no symbols that aren't one of those.

OpenStudy (anonymous):

\[\left(\frac{-4}{-4}+4\right)*4 \] thank you @KingGeorge .. This solution seems perfect.

OpenStudy (kinggeorge):

You're welcome.